OpenStudy (anonymous):

give any 2 example sums for each of the following: (a+b)whole square (a-b)whole square (a+b)(c+d) (a-b)(c+d) (a-b)(c-d)

OpenStudy (anonymous):

\[(a + b)²\] Example 1: (x + 3)² = (x + 3)(x + 3) = x² + 3x + 3x + 9 Distributive method = x² + 6x + 9 * Notice the middle term. Example 2: (x - 4)² = (x - 4)(x - 4) = x² - 4x - 4x + 16 Distributive method = x² - 8x + 16 * Again, notice the middle term.

OpenStudy (anonymous):

u said example

OpenStudy (anonymous):

(a+b)whole square a²+2ab +b²

OpenStudy (anonymous):

now find for (a - b)²

OpenStudy (anonymous):

got the answer

OpenStudy (anonymous):

@dharmikhokhani

OpenStudy (anonymous):

U WANT A QUESTION FOR THE FORMULA

OpenStudy (anonymous):

okay solve (2a + 5b)²

OpenStudy (anonymous):

@dharmikhokhani

OpenStudy (anonymous):

(4x + 3y)²

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

HEY WHAT YOU NEED HELP WITH

random231 (random231):

yeah exactly i cant understand the question!

OpenStudy (anonymous):

dude man bro, nobody understands the question.

OpenStudy (anonymous):

jamaican me crazy.

OpenStudy (anonymous):

... what are you even talking about

OpenStudy (anonymous):

u need to form questions in whch u can use the formulas above

OpenStudy (anonymous):

what formulas? you mean expressions?

OpenStudy (anonymous):

yup, rght

OpenStudy (anonymous):

for (a+b)^2, i guess you could have a problem like this: find the length of the hypotenuse